OpenStudy (anonymous):

The motion of a set of particles moving along the x-axis is governed by the differential equation dx/dt=t^3-x^3 where x(t) denotes the position at time t of the particle. a. If a particle is located at x=1 when t=2, what is its velocity at this time? b. Show that the acceleration of a particle is given by d^2x/dt^2=3t^2-3t^3x^2+3x^5 c. If a particle is located at x=2 when t=2.5, can it reach the location x=1 at any later time? (hint: t^3-X^3=(t-x)(t^2+xt+x^2)

OpenStudy (mathmale):

Part (a) is fairly straightforward. Any idea of what to do here, to determine the velocity of the particle when x=1 and t=2?

OpenStudy (anonymous):

I understand a. you just plug the values of x and t into the differential equation to get velocity: 2^3-1^3=7 Part b is the derivative of velocity differentiated implicitly with respect to t But I am lost on part c??

OpenStudy (mathmale):

I'll reply by posing a question for you: Is it possible to find x as a function of t through integration? If so, you could let x=1 and try solving for t. If there is a solution (or are solutions) to that, then you could say whether or not the particle ever returns to x=1.

OpenStudy (anonymous):

b is \[3t ^{2}-2x ^{2}(\frac{ dx }{ dt })\]

OpenStudy (mathmale):

If you start with the velocity and want a formula for distance, would you differentiate or integrate? Note: in Part c, you're given a hint: It says that when t=2.5, x=2.

OpenStudy (mathmale):

Is this a differential equations course, or just an advanced calculus course?

OpenStudy (anonymous):

I would integrate to get a formula for distance

OpenStudy (anonymous):

it is the beginning of a differential equations course

OpenStudy (anonymous):

i have to go to work right now but will be back later to try and solve this, thanks so far!!

OpenStudy (anonymous):

When I integrated both sides I got: \[\frac{ dx }{dt }=t ^{3}-x ^{3}\] \[x ^{3}dx=t ^{3}dt\] \[\int\limits x ^{3}dx=\int\limits t ^{3}dt\] \[\frac{ 1 }{ 4}x ^{4}=\frac{ 1 }{ 4 }t ^{4}+c\] Plugging in x=2 and t=5/2 I get c=-369/64 So if t is greater than 2.5 x can never be equal to 1

OpenStudy (mathmale):

First: dx/dt=t^3-x^3 = (t-x)(t^2+xt+x^2) (This hint is in the problem statement). Second: I see you've attempted to integrate dx/dt=t^3-x^3 as though it were a separable differential equation. I don't think it is one. If:\[\frac{ dx }{ dt }=t ^{3}-x ^{3}\] we could multiply both sides by dt to obtain:\[dx=t ^{3}dt - x^3dt\] but this does not convert to x^3 dx = t^3 dt, does it?

OpenStudy (mathmale):

I'm feeling conflicted here. On one had you say you're just starting a diff'l. eq'ns. course, but on the other hand, more sophisticated methods than ours are needed to find a solution to this d. e. in two variables. So I'm wondering whether there's a simpler, graphical approach, where you don't try to integrate dx/dt, but rather calculate the slope for a number of combos of x and t (such as x=2 and t=2.5). Have you done this before? I would think so, if you're beginning this d. e. course. If you go back to the equation for dx/dt, and let x=2 and t=2.5, you'd get dx/dt = (2.5)^3 - (2)^3, or 6.25 - 8, or -1.75. That's the same as -7/4. You could actually graph (2,2.5) and then draw in a short, downward sloping line to represent the slope of -7/4 at that point. If you were to do this again and again, getting closer to x = 1, and then sketch a rough graph that is always parallel to your little slope lines, you'd be able to see from the graph whether the graph would ever approach x=1 again. Hope this helps. I've been on my computer for hours and am no longer thinking very clearly.

OpenStudy (anonymous):

This problem is from the 3rd section of chapter 1 in differential equations. The section covers direction fields.

OpenStudy (anonymous):

I'm pretty confused on part c of this problem. When I plug in the values t=2.5 and x=2 I get dx/dt=(2.5)^3-2^3=7.625

OpenStudy (anonymous):